OpenStudy (anonymous):

-5 5/8 divided by (-4 7/12)=

OpenStudy (campbell_st):

change both mixed numbers improper fractions. \[-\frac{5 \times 8 + 5}{8} = -\frac{45}{8}\] and \[-\frac{4 \times 12 + 7}{12} = -\frac{55}{12}\] now the problem is \[-\frac{45}{8} \div -\frac{55}{12}\] the rule for dividing fractions is, flip the 2nd fraction and multiply \[-\frac{45}{8} \times -\frac{12}{55} = \] you also need to know, when multiplying that 2 negatives will create a positive... so your answer will be positive hope it helps

OpenStudy (anonymous):

thanks